# A Simple-but-effective Baseline for Training-free Class-Agnostic Counting

### v1

@@ 1 Introduction
Recently, the Segment Anything Model (SAM) [9] has garnered substantial attention across various research domains due to its remarkable ability to generate high-quality object masks for any given image. This versatility has led to its application in numerous research areas, as evidenced by multiple studies  [27, 15, 25, 21]. In the context of CAC, researchers have started adopting SAM for producing instance-level segmentation masks corresponding to referenced objects. By counting the number of instance masks, they can determine the final count of objects. Thanks to SAM’s robust instance segmentation capabilities, this counting-via-instance-segmentation approach can function effectively even when utilizing the off-the-shelf SAM model without any additional training. Training-free methods hold appeal for CAC since they eliminate the need for a dedicated training dataset and alleviate the challenges associated with collecting extra training data. Nevertheless, despite their promise, current training-free CAC methods still lag behind their training-based counterparts in terms of performance, e.g., as demonstrated in Figure 1, the current top training-free method falls short of the best training-based approaches by 84.89% and 105.82% on the FSC-147 [19] and CARPK [8] datasets, respectively.
@@
Through our rigorous experimental evaluation, we have demonstrated the remarkable performance of our proposed methods and the effectiveness of the four key components. Significantly, we have shown, for the first time, that a Class-Agnostic Counting (CAC) model, even without any training, can achieve performance on par with that of training-based CAC approaches.
@@ 3.1 Object-prior Point Prompt Matters
In this study, we use the classic superpixel algorithm named Simple Linear Iterative Clustering (SLIC) which can generate compact superpixels with a low computational overhead [2]. Given the initial cluster number $K$, SLIC can directly output superpixel centers $\mathcal{C}=\{(x_{i},y_{i})\}_{i=1}^{K}$ which can then be fed into SAM together with the provided reference boxes (or points) $\mathcal{R}$ batch by batch for mask proposal generation
@@
where $\mathbb{E}$ and $\mathbb{D}$ are encoder and decoder of SAM respectively. $I$ is the input image. $\mathcal{M}_{\text{ref}}$ denotes the generated masks for corresponding referenced prompts and $\mathcal{M}_{\text{proposal}}$ means all potential mask proposals  [^3] The biggest mask for background segmentation and masks for referenced objects are filtered from $\mathcal{M}_{\text{proposal}}$ for simplicity, i.e., $\mathcal{M}_{\text{proposal}}\cap\mathcal{M}_{\text{ref}}=\emptyset$..
@@ 3.2 Semantic-rich Feature Matters
Next, an overall segmentation for interested objects can be obtained by selecting mask proposals whose similarity score is higher than a predefined threshold $\theta$. However, searching for an optimal $\theta$ can be hard due to the feature representations of SAM lacking clear semantic-related knowledge [16, 10]. As empirical analysis presented in Figure 4, many objects of different categories from the reference object may have a quite close similarity score to that of objects in the same category, thus it is inevitable to output poor counting results with noisy segment generation. For instance, the SAM feature struggles to differentiate between apples and other messy objects in row 1, as well as children’s hair and feet in row 2 of Figure 4.
@@
In order to obtain a plausible counting accuracy, using semantic-rich features for mask proposal selection becomes indispensable. Recently, many vision foundation models with rich semantic knowledge have been made accessible to the public, such as DINO [3, 17]. Combining with the semantic-rich feature representation, the final counting results can be obtained
@@ 3.3 Multi-scale Segmentation Helps
Training-free Reference Format FSC-147 CARPK
@@
GMN [14] [ACCV 19’] No Box 26.52 124.57 9.90 -
FamNet [19] [CVPR 21’] No Box 22.08 99.54 18.19 33.66
CFOCNet+ [28] [WACV 21’] No Box 22.10 112.71 - -
BMNet+ [22] [CVPR 22’] No Box 14.62 91.83 5.76 7.83
SAFECount [29] [WACV 23’] No Box 14.32 85.54 5.33 7.04
LOCA [6] [ICCV 23’] No Box 10.79 56.97 9.97 12.51
SAM Baseline [23] Yes N.A. 42.48 137.50 16.97 20.57
Count-Anything [16] Yes Box 27.97 131.24 - -
TFOC [23] [WACV 24’] Yes Box 19.95 132.16 10.97 14.24
Ours Yes Box 12.26 56.33 4.39 5.70
TFOC [23][WACV 24’] Yes Point 20.10 132.83 11.01 14.34
Ours Yes Point 12.47 49.97 4.39 5.70
@@ 4.1 Experimental Setup
All experiments in this work are conducted on one NVIDIA GeForce RTX 3090 GPU and all other experimental settings are identical to the TFOC [23] approach for a fair comparison.
@@ 4.2 Comparison with State-of-the-Arts
Qualitative Analysis: Figure 6 displayed the visualization results of various counting methods in different scenarios and our approach can always more accurately identify objects of interest in various scenes. Specifically, for scenes in the first three rows, both of the compared training-based and training-free methods mistakenly identified areas that do not belong to the interested object, such as the teacup handle and the spoon in row 1, the crowded leaves in row 2 and the table legs and the top edge of the chair back in row 3. While our approach can successfully and accurately segment out the interested object and produce a more accurate counting result. This will be attributed to the utilization of semantic-rich feature representations for SAM’s mask proposal selection in our approach. Furthermore, for more challenging scenes in the last two rows, the density maps generated by the training-based BMNet+ method are vague and exhibit substantial deviations in counting outcomes. The TFOC method, on the other hand, either erroneously includes extensive background areas in its object count or overlooks numerous objects in crowded scenes. Yet, our method yields accurate segmentation masks for individual objects due to its multi-scale mechanism, resulting in output counting numbers that more closely align with the ground truth.
@@ 5 Ablation study
Table 2: Performance with different components. SP: superpixel, TPU: transductive prototype updating, MS: multi-scale.
@@
Effectiveness of different components: Table 2 showcases the impact of each component in our method. Specifically, although SAM with superpixel can achieve higher interested-object recall (shown in Figure 2), due to the lack of semantic differentiation ability in SAM features (shown in Figure 4), simply replacing default point grid with superpixel does not necessarily bring performance gains, as the results shown in the first two rows of Table 2. The next two rows show that using DINOv2 feature representation for SAM’s mask proposal selection significantly reduces the final counting error and the combing of superpixel can further bring performance gains. The subsequent two rows highlight the individual benefits of transductive prototype updating and the multi-scale mechanism on our model’s effectiveness. When these four components are combined, they lead to the most impressive counting results.
@@
Various semantic-rich vision models: Table 4 displays our method’s performance across different vision models, indicating that superior vision models yield improved counting accuracy.
@@
Figure 7: Performance with various threshold $\theta$ and $\delta$.
@@
Threshold sensitivity: Our method involves two hyperparameters: a similarity threshold $\theta$ for proposal selection and $\delta$ for transductive prototype updating. Figure 7 illustrates the robustness of our approach to variations in these hyperparameters.
@@ 6 Conclusion
[…]
Supplementary Material
In this supplementary material, we provide more details to complement the manuscript, including visualization of generated mask proposal in Sec. 7, CAC with text references in Sec. 8 and qualitative results on CARPK dataset in Sec. 9.
7 Visualization of Generated Mask Proposal
Figure 8: Visualization of mask proposals generated by SAM with various point prompts. In this example, the hot-air balloon is the user-interested object and the bottom row presents the recall of this object. The red boxes display mask generation for a very fine object (the street lamp).
In the main paper, we have presented the effectiveness of the use of superpixel on the quality of mask proposals generated by SAM in Figure 3. In this section, we further illustrate the visualization of generated mask proposals in Figure 8. It clearly shows that SAM with superpixel can achieve a high recall rate for the interested object without demanding of denser grid of points. Meanwhile, we find that SAM with superpixel can also generate mask proposals for very thin objects thanks to the object-prior prompt, such as the street lamp marked in the red box. Nonetheless, employing SAM with a standard grid of points might not consistently achieve segmentation for such objects.
8 CAC with Text References
In the main paper, the efficiency of our method using box and point references has been confirmed. In this section, we extend our examination to include the use of text references in testing our approach. For a fair comparison, we adopt the method from TFOC [23] and employ the CLIP-Surgery model to derive coarse bounding boxes for the target object using text references. Then our approach can use these coarse bounding boxes for object counting. As presented in the following Table 6, our approach is significantly superior to the compared methods.
Training-free FSC-147
MAE $\downarrow$ RMSE $\downarrow$
SAM Baseline [23] Yes 42.48 137.50
TFOC [23] (vanilla) Yes 32.86 142.89
TFOC [23] Yes 24.79 137.15
Ours Yes 14.88 56.91
Table 6: Effect of our approach with text prompt on FSC-147.
9 Qualitative Results on CARPK
In this section, we further provide some qualitative visualization for the CARPK dataset. As shown in Figure 9, our method can generate precise masks for each car instance and achieve accurate counting results.
Figure 9: Qualitative results on the CARPK dataset are displayed. Counting values are noted in the bottom-left corner. Best viewed by zooming in.

### v2

@@ 1 Introduction
Recently, the Segment Anything Model (SAM) [10] has garnered substantial attention across various research domains due to its remarkable ability to generate high-quality object masks for any given image. This versatility has led to its application in numerous research areas, as evidenced by multiple studies  [28, 16, 26, 22]. In the context of CAC, researchers have started adopting SAM for producing instance-level segmentation masks corresponding to referenced objects. By counting the number of instance masks, they can determine the final count of objects.
Thanks to SAM’s robust instance segmentation capabilities, this counting-via-instance-segmentation approach can function effectively even when utilizing the off-the-shelf SAM model without any additional training. Training-free methods hold appeal for CAC since they eliminate the need for a dedicated training dataset and alleviate the challenges associated with collecting extra training data. Nevertheless, despite their promise, current training-free CAC methods still lag behind their training-based counterparts in terms of performance, e.g., as demonstrated in Figure 1, the current top training-free method falls short of the best training-based approaches by 84.89% and 105.82% on the FSC-147 [20] and CARPK [9] datasets, respectively.
@@
Our rigorous experiments demonstrated the remarkable performance of our proposed methods and the effectiveness of the four key components. Significantly, we show for the first time that a Class-Agnostic Counting (CAC) model without any training can achieve performance on par with that of training-based CAC approaches.
@@ 3.1 Object-prior Point Prompt Matters
In this work, a superpixel algorithm named Simple Linear Iterative Clustering (SLIC) is used to generates compact superpixels with a low computational overhead [1].
@@
Given the initial cluster number $K$, SLIC can directly output superpixel centers $\mathcal{C}=\{(x_{i},y_{i})\}_{i=1}^{K}$. These centers, along with the provided reference boxes (or points) $\mathcal{R}$, can then be fed into SAM batch by batch for mask proposal generation
@@
where $\mathbb{E}$ and $\mathbb{D}$ are encoder and decoder of SAM respectively. $I$ is the input image. $\mathcal{M}_{\text{ref}}$ denotes the generated masks for corresponding referenced prompts (reference boxes or points) and $\mathcal{M}_{\text{proposal}}$ means all potential mask proposals 333 The largest mask generated by SAM, which typically corresponds to the background, and masks for referenced objects are filtered from $\mathcal{M}_{\text{proposal}}$ for simplicity, i.e., $\mathcal{M}_{\text{proposal}}\cap\mathcal{M}_{\text{ref}}=\emptyset$..
@@ 3.2 Semantic-rich Feature Matters
Next, an overall segmentation for interested objects can be obtained by selecting mask proposals whose similarity score is higher than a predefined threshold $\theta$. However, searching for an optimal $\theta$ can be hard due to the feature representations of SAM lacking clear semantic-related knowledge [17, 11]. As empirical analysis presented in Figure 4, many objects of different categories from the reference object may have a quite close similarity score to those objects of the same category, thus it is inevitable to output poor counting results with noisy segment generation. For instance, the SAM feature struggles to differentiate between apples and other messy objects in row 1, as well as children’s hair and feet in row 2 of Figure 4. Moreover, our empirical findings (the first two rows in Table 3) indicate that the enhanced recall of referenced objects achieved through the proposed superpixel method does not improve the final counting performance when the SAM feature representation is directly employed for object discrimination. Therefore, to obtain a plausible counting accuracy, using semantic-rich features for mask proposal selection becomes indispensable.
@@
Recently, many vision foundation models with rich semantic knowledge have been made accessible to the public, such as DINO [3, 18].
Combining with the semantic-rich feature representation, the final count can be obtained
@@ 3.3 Multi-scale Segmentation Helps
Training Reference Format FSC-147 CARPK
@@
GMN [15] [ACCV 19’] Yes Box 26.52 124.57 9.90 -
FamNet [20] [CVPR 21’] Yes Box 22.08 99.54 18.19 33.66
CFOCNet+ [29] [WACV 21’] Yes Box 22.10 112.71 - -
BMNet+ [23] [CVPR 22’] Yes Box 14.62 91.83 5.76 7.83
SAFECount [30] [WACV 23’] Yes Box 14.32 85.54 5.33 7.04
PseCo [7] [CVPR 24’] Yes Box 13.05 112.86 - -
LOCA [6] [ICCV 23’] Yes Box 10.79 56.97 9.97 12.51
SAM Baseline [24] No N.A. 42.48 137.50 16.97 20.57
Count-Anything [17] No Box 27.97 131.24 - -
TFOC [24] [WACV 24’] No Box 19.95 132.16 10.97 14.24
Ours No Box 12.26 56.33 4.39 5.70
TFOC [24][WACV 24’] No Point 20.10 132.83 11.01 14.34
Ours No Point 12.47 49.97 4.39 5.70
TFOC [24][WACV 24’] No Text 24.79 137.15 - -
Ours No Text 23.59 113.60 - -
@@ 4.1 Experimental Setup
All experiments in this work are conducted on one NVIDIA GeForce RTX 3090 GPU and all other experimental settings (including the text prompt) are identical to the TFOC [24] approach for a fair comparison.
@@ 4.2 Comparison with State-of-the-Arts
Training FSC-147
MAE $\downarrow$ RMSE $\downarrow$
FamNet [20] [CVPR 21’] Yes 26.76 110.95
BMNet+ [23] [CVPR 22’] Yes 16.89 96.65
LOCA [6] [ICCV 23’] Yes 12.53 75.32
TFOC [24] [WACV 24’] No 22.96 137.27
Ours No 15.15 69.16
Table 2: Evaluation on one-shot setting on FSC-147 test set.
@@
We also evaluate our pipeline on FSC-147 test set with a one-shot counting scenario. Table 2 indicates that our approach is robust on a one-shot counting scenario, which is consistent with the results of the three-shot above.
And for text prompts, we followed TFOC [24] and employ the CLIP-Surgery model to derive coarse bounding boxes for the target object using text references. Then our approach use these coarse bounding boxes for object counting. As presented in last rows of the Table 1, our approach is sightly superior to the compared methods. The main reason is that the bounding boxes generated from TFOC [24] are not accurate. Technically, great improvement will be observed if high quality bounding boxes are given.
Qualitative Analysis: Figure 6 displayed the visualization results of various counting methods in different scenarios and our approach can always more accurately identify objects of interest in various scenes. Specifically, for scenes in the first three rows, both of the compared training-based and training-free methods mistakenly identified areas that do not belong to the interested object, such as the teacup handle and the spoon in row 1, the crowded leaves in row 2 and the table legs and the top edge of the chair back in row 3. While our approach can successfully and accurately segment out the interested object and produce a more accurate counting result. This will be attributed to the utilization of semantic-rich feature representations for SAM’s mask proposal selection in our approach. Furthermore, for more challenging scenes (the last two rows), the density maps generated by the training-based BMNet+ method are vague and exhibit substantial deviations in counting outcomes. The TFOC method either erroneously includes extensive background areas in its object count or overlooks numerous objects in crowded scenes. Yet, our method yields accurate segmentation masks for individual objects due to its multi-scale mechanism, resulting in output counting numbers that more closely align with the ground truth.
@@ 5 Ablation study
Table 3: Performance with different components on FSC-147. SP: superpixel, TPU: transductive prototype updating, MS: multi-scale.
Effectiveness of different components: Table 3 showcases the impact of each component in our method in both dataset. Specifically, although SAM with superpixel can achieve higher interested-object recall (shown in Figure 2), due to the lack of semantic differentiation ability in SAM features (shown in Figure 4), simply replacing default point grid with superpixel does not necessarily bring performance gains, as the results shown in the first two rows of Table 3. The next two rows show that using DINOv2 feature representation for SAM’s mask proposal selection significantly reduces the final counting error and the combing of superpixel can further bring performance gains. The subsequent two rows of Table 3 highlight the individual benefits of transductive prototype updating and the multi-scale mechanism on our model’s effectiveness. When these four components are combined, they lead to the most impressive counting results, especially in the RMSE metric..
number of K MAE $\downarrow$ RMSE $\downarrow$ avg. time/image
50*50 5.09 6.66 5.26s
60*60 4.46 5.82 6.55s
64*64 4.39 5.70 7.04s
Table 4: Performance of our method with various K on CARPK.
Different number of initialized points: Table 4 shows that with 50*50 points, our method surpasses the previous SOTAs (see Table 1) in reasonable time. Increasing to 64*64 points further improves the MAE by 18%, while the computational time remains within seconds, demonstrating our method’s ability to achieve significant performance gains with manageable time costs as the point density increases.
@@
Various semantic-rich vision models: Table 6 displays our method’s performance across different SOTA foundation vision models, indicating that superior vision models yield improved counting accuracy. For CLIP, only the vision encoder is involved.